P(A) = 0.30
P(B) = 0.40
P(C) = 0.10
P(AorD) = 0.50
P(A and D ) = 0.04
P(A and C) = 0.03
P(B/A) = 0.60
Determine :
P(AandB)
P(A/B)
P(D)
thank you

Answer by stanbon(48568)

P(A) = 0.30
P(B) = 0.40
P(C) = 0.10
P(A or D) = 0.50
P(A and D ) = 0.04
P(A and C) = 0.03
P(B|A) = 0.60
--------------------
Determine :
P(A and B) = P(A)*P(B|A) = 0.3*0.6 = 0.18
------------
P(A|B) = P(A and B)/P(B) = 0.18/0.4 = 0.45
--------------
P(D) = 1 - 0.3-0.4-0.1 = 0.2
===============
Cheers,
Stan H.